OpenStudy (anonymous):

What is the solution of 4x - 2 = 4x - 2 and why?

OpenStudy (anonymous):

No solution, everything cancels out

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yup 4x-4x=-2+2 =0 every thing gets cancelled